# ELOQ: Resources for Enhancing LLM Detection of Out-of-Scope Questions

### v3

@@ Abstract.
Large Language Models (LLMs) are widely used in Conversational AI systems to generate responses to user inquiries. However, many natural questions lack well-defined answers. While existing studies primarily focus on question types such as false premises, they often overlook out-of-scope questions, where the provided document is semantically highly similar to the query but does not contain the required answer. In this paper, we propose a guided hallucination-based method to efficiently generate a diverse set of out-of-scope questions from a given document corpus. We then evaluate multiple LLMs based on their effectiveness in confusion detection and appropriate response generation. Furthermore, we introduce an improved method for detecting such out-of-scope questions, enhancing the reliability of LLM-based question-answering systems.
Large Language Models, Question Answering, Out-of-Scope Question
*Both authors contributed equally to this work.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processing
@@ 1. Introduction
Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents (Lewis et al., 2020; Guu et al., 2020; Gao et al., 2024). Given an inquiry, a RAG system retrieves a relevant document from a curated knowledge base and provides it to a large language model (LLM) that generates a response. RAG performance is evaluated on dimensions such as response accuracy, faithfulness, completeness, answer and context relevance, and style (Es et al., 2024; Saad-Falcon et al., 2024; Liu et al., 2024; Zheng et al., 2023). Errors may stem from a document-to-question mismatch, hallucinated parts of the response, or misunderstanding of the document or the question.
Many user questions have no direct answer: some rely on a false premise, others are ambiguous in their intent, and some are out-of-scope for the knowledge base. Prior studies indicate that about 25% of natural questions contain false assumptions (Yu et al., 2023) and over 50% are ambiguous (Min et al., 2020a). Additionally, out-of-scope errors have been identified as the leading cause of enterprise AI assistants generating responses that appear convincing but are, in fact, incorrect (Maharaj et al., 2024). A well-designed RAG system should implement an exception-handling mechanism that detect confusing questions and responds by clarifying the confusion, retrieving additional documents, or seeking assistance.
Dataset Source Category Cutoff knowledge Annotation Method
@@
ELOQ News Out-of-scope Human+LLM
Table 1. ELOQ vs Existing Datasets
Enabling “deconfusion” in a RAG system involves three challenges: (1) creating a diverse dataset of confusing questions, (2) training a classifier to detect confusion, and (3) developing a response generator for each confusion type. As shown in Table 1, much prior work focuses on false-premise questions that contradict pretraining knowledge, but often neglects out-of-scope queries for which no answers can be found in retrieved documents. Out-of-scope questions can mislead LLMs into generating hallucinated answers that appear correct, but are incorrect, which represents the most dangerous type of error (Maharaj et al., 2024). Our focus is to mitigate this type of confusion.
This paper proposes a framework for automatically generating out-of-scope questions and evaluates its effectiveness in training classifiers to detect out-of-scope situations. While queries in MS MARCO (Nguyen et al., 2016) may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cutoff. ELOQ fills this gap by providing a diverse benchmark for out-of-scope questions, leveraging LLM-assisted generation to reduce human annotation effort. Our contributions are as follows:
  • We propose and implement an efficient method for generating synthetic out-of-scope questions from a given set of documents.
  • We demonstrate the utility of our synthetic data in training out-of-scope exception detectors.
  • We conduct a comparative evaluation of multiple LLMs as RAG agents to measure their confusion detection accuracy.
  • Our code  [^1] https://github.com/zhiyuanpeng/ELOQ.git and dataset  [^2] https://huggingface.co/datasets/zhiyuanpeng/ELOQ are made publicly available.
@@ 2.1. Benchmark Datasets
Recent work on LLM responses to confusing questions primarily focuses on stand-alone questions answered using the LLM’s pretraining knowledge, without a supporting context document. These questions may be naturally collected from online sources (Li et al., 2024b; Yu et al., 2023), written by human annotators (Hu et al., 2023), or synthetically generated (Zhu et al., 2024; Yuan et al., 2024). Released benchmark datasets with confusing questions include KG-FPQ (Zhu et al., 2024), FalseQA (Hu et al., 2023), FLUB (Li et al., 2024b), and CREPE (Yu et al., 2023). Earlier research on ambiguous open-domain questions (Min et al., 2020b) introduced AmbigNQ, a dataset of naturally ambiguous questions. Table 1 presents a comparison of these benchmark datasets.
@@ 2.2. Question Generation
(Zhu et al., 2024; Yuan et al., 2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of (subject, relation, object) using Wikipedia data, replacing the object with a similar but incorrect one, and then filling in a template to generate a false premise question. Content-grounded (non-confusing) synthetic data generation has been studied in (Yehudai et al., 2024; Wang et al., 2023a; Lee et al., 2023). Given the grounding passages, these methods generate data entries using few-shot prompting with task-specific examples, followed by filtering to ensure data quality, faithfulness, and diversity.
@@ 2.3. Mitigation
Mitigation is commonly achieved through in-context learning  (Zhu et al., 2024; Li et al., 2024b), chain-of-thought (CoT)  (Wei et al., 2022) and fine-tuning. Experiments show that fine-tuning outperforms both in-context learning and CoT reasoning (Hu et al., 2023). CoT reasoning leads to inconsistent improvements, whereas in-context learning improves performance as the number of examples increase (Li et al., 2024b). Yuan et al. (2024) identifies approximately 1% of attention heads as responsible for confusion caused by false-premise questions and demonstrates a 20% performance gain by constraining these heads.  Kim et al. (2021) retrieves a Wikipedia page for each question and responds with the identified unverifiable premises if the page does not entail them.
@@ 3. Methods
We developed the ELOQ data generator to evaluate and mitigate LLM confusion when responding to out-of-scope questions for a given document. We assume that the user has limited knowledge of the document, resulting in asking questions that appear relevant but are actually out-of-scope and cannot be answered based on the document’s content. In this step, given a standard prompt $p$, an out-of-scope question $q$, and a relevant document $d$, the LLMr generates a response $r$. Ideally, $r$ should clarify the confusing part of the question – a process we refer to as defusion (or de-confusion) – rather than attempting to answer the question directly and risking hallucination.
@@
Attr Gold Silver
# of documents 45 1955
# of in-scope questions 103 8,949
# of out-of-scope questions 113 10,105
average words per document 280 276
average words per in-scope question 16 16
average words per out-of-scope question 16 16
Table 2. Statistics of ELOQ. Gold is sampled from crawled 2000 news and annotated by annotators (Section 3.4).
3.1. Data Collection
We focus on the scenario where the domain knowledge is maintained separately from the LLM in a document database. Hence, we prefer documents that are novel to the evaluated LLMs, such as news articles published after all pretraining cutoff dates, ensuring that LLMs cannot reproduce claims or answer questions correctly by pure hallucination. Instead, they must generate responses grounded in their understanding of the news content. Therefore, using the Newscatcher (Bugara et al., 2020), we collected 200 news articles published between Jan 1, 2024, and September 30, 2024, for each topic333sport, business, science, food, politics, travel, entertainment, music, news, tech. We require each document to be short enough to fit in the LLM prompt along with instructions and examples but long enough to make at least 4 to 10 separate claims (i.e. a few paragraphs long). Thus, we select news articles with more than 150 words and, for each, sequentially extract sentences from the beginning until the word count exceeds 300. Finally, we collected 2,000 news and sampled a subset for human annotation. We represent the human-annotated data as Gold and the remaining data as Silver. Table 2 presents the dataset statistics.
3.2. Question Generation
3.2.1. Claim Extraction
$S_{j}=\{c_{i}\mid i\mod 3=j-1\},\quad\text{for }j=1,2,3.$
3.2.2. Hallucination Injection
The second step is to generate hallucinated claims that are similar to the factual claims generated in Section 3.2.1 but contain altered or missing information. The process follows an iterative masking and recovery approach over multiple rounds (e.g. 3). For each round:
  • For all claims in one subset $S_{j}$, we replace them with placeholders “(missing)”.
  • The modified claims list (with missing claims) is then provided to LLMq without access to $d$. The model is prompted to recover the missing claims based on the remaining claims in the list.
  • The recovered claims are inserted back into the claims list, replacing the placeholders.
  • This process is repeated for all subsets $S_{1},…,S_{k}$ across several rounds.
Finally, we discard claims that are explicitly supported by $d$. This ensures that only truly out-of-scope claims remain in our dataset.
3.2.3. Question Generation
In addition to generating out-of-scope questions, we also prompt LLMq to generate in-scope questions for comparison. The complete question generation process, along with its prompts, is outlined in Algorithm 1.
Group Name Cohen’s Kappa Annotator Acc Group’s Agree Acc Ground Truth Acc
1   — 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
  — 87.50 94.59
2   — 0.7508 0.8187 93.06 97.14 98.41 96.77
  — 91.67 90.00
3   — 0.8615 0.8333 91.67 93.75 98.51 100.00
  — 98.61 100.00
Table 3. Agreement scores and the performance of our method on the ELOQ-Gold.
@@ 4. Experimental Results
Table 5. Evaluation of LLMs’ accuracy in defusing out-of-scope questions from the ELOQ-Silver dataset across diverse news topics. For each LLM, the underscored value and stared () value are the minimum and maximum values, respectively, across all the topics within the same prompting method. Bold values are the maximum values for each topic across all the LLMs within the same prompting method. “entm” is the abbreviation of “entertainment”.
@@ 4.2. Out-of-Scope Detection
The main challenge in preventing an LLM from hallucinating or generating incorrect answers to out-of-scope questions is detecting such scenarios. Once we can accurately identify these cases, we can handle them easily by using a prompt and in-context examples specifically designed for out-of-scope questions. We use two methods for out-of-scope detection. First is a prompt-based method where we simply prompt the LLM to do such detection (Appendix B.3), denoted as “Direct Generate” in Figure 2. Second is a classification method where we train a binary classifier on ELOQ-Silver using the embeddings of an appended unused-token as features, denoted as “Unused-token Classification” in Figure 2.
@@ 4.2.1. Unused-token Classification
The out-of-scope detection problem can be framed simply as a binary classification task, where we determine if a given query $q$ is out-of-scope or not for a given document $d$. The binary label is $y\in\{0,1\}$ where $y=1$ indicates the query is relevant to the document and $y=0$ indicates the query is out-of-scope. The objective is to learn a classification function: $f:(q,d)\mapsto p(y=1\mid q,d)$, which estimates the likelihood of the query-document pair to be in-scope. We would like to test whether the LLM’s internal features contain enough information to classify a query-document pair effectively to simulate the effect that the LLM “realizes” the out-of-scope scenario before generating a response. To achieve this, we use the sequence representation of the entire input – denoted as $x=\text{concat}(I,q,d,)$ where $I$ is the instruction prompt provided to the LLM (e.g. “Can this document answer the question?”). We extract the sequence representation by appending a reserved unused-token to the end of the input sequence and extracting its representation from the final hidden state of the last transformer layer. The rationale behind using unused-token is that, since LLM developers reserve these tokens for potential future use in specialized training tasks, they are untrained and uninfluenced by any prior data. Thus, the unused-token does not have any predefined meanings and can be viewed as a blank buffer, aggregating interactions among all the input tokens and capturing an uninfluenced representation of the sequence. We also tried using the <eos> token, which yielded similar results.
The classifier is a two-layer MLP that takes the unused-token representation as input and is optimized against the binary label $y$. This approach is commonly known as “probing” (Ye et al., 2024), where the LLM weights are frozen, and the classifier can be viewed as an external probe to assess the information contained in the model’s internal representations. The performance reported in Figure 2 was tested on ELOQ-Gold, the human-labeled data. For more details about the training, please refer to Appendix D. We also provide a “Sentence BERT NLI” baseline where we adopt NLI style training using the off-the-shelf  [^4] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 model from HuggingFace. Specifically, we follow (Reimers and Gurevych, 2019) to extract document and question [CLS] embeddings $e_{d}$ and $e_{q}$ to use $[e_{q};e_{q};|e_{d}-e_{q}|]$ as the input features.

### v4

@@ Abstract.
Retrieval-augmented generation (RAG) has become integral to large language models (LLMs), particularly for conversational AI systems where user questions may reference knowledge beyond the LLMs’ training cutoff. However, many natural user questions lack well-defined answers, either due to limited domain knowledge or because the retrieval system returns documents that are relevant in appearance but uninformative in content. In such cases, LLMs often produce hallucinated answers without flagging them. While recent work has largely focused on questions with false premises, we study out-of-scope questions, where the retrieved document appears semantically similar to the question but lacks the necessary information to answer it. In this paper, we propose a guided hallucination-based approach ELOQ  [^1] https://github.com/zhiyuanpeng/ELOQ.git, to automatically generate a diverse set of out-of-scope questions from post-cutoff documents, followed by human verification to ensure quality. We use this dataset to evaluate several LLMs on their ability to detect out-of-scope questions and generate appropriate responses. Finally, we introduce an improved detection method that enhances the reliability of LLM-based question-answering systems in handling out-of-scope questions.
Large Language Models, Question Answering, Out-of-Scope Question, Retrieval Augmented Generation
ccs: Computing methodologies Natural language processingccs: Information systems Question answering
@@ 1. Introduction
Retrieval Augmented Generation (RAG) has become a standard approach to building context-grounded conversational AI agents (Lewis et al., 2020; Guu et al., 2020; Gao et al., 2024). Given an inquiry, a RAG system retrieves a relevant document from a curated knowledge base and presents it to a large language model (LLM), which generates a response. RAG performance is evaluated on dimensions such as response accuracy, faithfulness, completeness, answer and context relevance, and style (Es et al., 2024; Saad-Falcon et al., 2024; Liu et al., 2024; Zheng et al., 2023). Errors may stem from a document-to-question mismatch, hallucinated parts of the response, or misunderstanding of the document or the question.
Many user questions have no direct answer: some rely on a false premise, others are ambiguous in their intent, and some are out-of-scope for the knowledge base. Prior studies indicate that about 25% of natural questions contain false assumptions (Yu et al., 2023) and over 50% are ambiguous (Min et al., 2020a). Additionally, out-of-scope errors have been identified as the leading cause of enterprise AI assistants generating responses that appear convincing but are, in fact, incorrect (Maharaj et al., 2024). A well-designed RAG system should implement an exception handling mechanism that detects confusing questions and responds by clarifying the confusion, retrieving additional documents, or seeking assistance.
Dataset Source Category Cutoff Annotation Method
SQuAD 2.0 (Rajpurkar et al., 2018) Wiki Unanswerable Human
SQuAD2-CR (Lee et al., 2020) Wiki Unanswerable Human
BoolQ3L (Sulem et al., 2022) Wiki Unanswerable Algoritm+Human
CAsT-answerability (Lajewska and Balog, 2024) Wiki Unanswerable Algoritm+Human
@@
ELOQ News Out-of-scope LLM+Human
Table 1. ELOQ vs Existing Datasets. “Cutoff” represents whether the data is beyond the LLMs’ knowledge or not.
Enabling “deconfusion” in a RAG system involves three challenges: (1) creating a diverse dataset of confusing questions, (2) training a classifier to detect confusion, and (3) developing a response generator for each confusion type. As shown in Table 1, prior studies create confusing questions with different categories, such as multi-answer (Min et al., 2020a), false premise (Yu et al., 2023; Hu et al., 2023; Zhu et al., 2024; Yuan et al., 2024), and cunning texts (Li et al., 2024a). For instance, “How many sacks does clay matthews have in his career” is a multi-answer question as “clay matthews” may refer to “Clay Matthews Jr.” or “Clay Matthews III”, “Where in Liverpool did John Lennon die” is a false premise question as John Lennon died in New York city and “What should I do if I forget which ATM machine I deposited my money in?” is a cunning text. In contrast to these questions, out-of-scope questions look relevant to the document, but there is no answer within the document and thus can mislead LLMs into generating hallucinated answers that seem correct but are actually incorrect, representing the most dangerous type of error (Maharaj et al., 2024). For instance, “How does Justin Jefferson’s exceptional catching ability influence the Vikings’ decision to select him in the first round of the 2020 NFL Draft despite his lack of speed?” is an out-of-scope question as it asks about “Vikings’ decision to select Jefferson in the first round of the 2020 NFL Draft” which was never discussed in the paired news (Appendix D). Our work focuses on mitigating this type of confusion.
Existing studies (Rajpurkar et al., 2018; Lee et al., 2020; Sulem et al., 2022; Lajewska and Balog, 2024) on unanswerable questions of which out-of-scope questions are a subset, have typically created the dataset either through manual annotation by trained human annotators (Lee et al., 2020; Hu et al., 2023) or via simple automatic generation algorithms, as seen in BoolQ3L (Sulem et al., 2022) and CAsT-answerability (Lajewska and Balog, 2024). These methods often lack a balance between efficiency and quality, as human-annotated questions are of high quality but low efficiency, and automatically generated questions are of high efficiency but too simple to be distinguished to some degree. Recently, LLMs have demonstrated strong generation and following instruction abilities, and thus KG-FPQ (Zhu et al., 2024) and FAITH (Yuan et al., 2024) utilize LLMs to generate high-quality false premise questions, mitigating the balance of efficiency and quality issues. However, all these datasets are usually constrained to knowledge before the LLMs’ training cutoff date, which limits their effectiveness in evaluating real-world RAG systems that must retrieve up-to-date information likely beyond the model’s training data to answer user questions.
To address these challenges above, we propose a framework for automatically generating out-of-scope questions and evaluate its effectiveness in training classifiers to detect out-of-scope situations. While questions in existing datasets such as MS MARCO (Nguyen et al., 2016), may serve as out-of-scope questions for the corresponding hard negative passages, they fail beyond the LLM’s knowledge cutoff. ELOQ fills this gap by providing a diverse benchmark for out-of-scope questions, leveraging LLM-assisted generation to reduce human annotation effort. Our contributions are as follows:
  • We propose a framework for generating synthetic out-of-scope questions from a given corpus.
  • We demonstrate the utility of our synthetic data in training out-of-scope question detectors.
  • We conduct a comparative evaluation of multiple LLMs as RAG agents to measure their out-of-scope question detection accuracy.
@@ 2.1. Benchmark Datasets
Recent work on LLM responses to confusing questions primarily focuses on stand-alone questions answered using the LLM’s pretraining knowledge, without a supporting context document. These questions may be naturally collected from online sources (Li et al., 2024b; Yu et al., 2023), written by human annotators (Hu et al., 2023), or synthetically generated (Zhu et al., 2024; Yuan et al., 2024). Earlier research on ambiguous open-domain questions introduced AmbigQA (Min et al., 2020b), a dataset of naturally ambiguous questions. More relevant benchmarks are unanswerable questions (Rajpurkar et al., 2018; Lee et al., 2020; Sulem et al., 2022; Lajewska and Balog, 2024) of which out-of-scope questions are a subset. SQuAD 2.0 (Rajpurkar et al., 2018) is created by asking crowdworkers to write unanswerable questions similar to the answerable question given a paragraph. SQuAD2-CR (Lee et al., 2020) tags SQuAD 2.0 (Rajpurkar et al., 2018) with answerable reasons. CAsT-answerability (Lajewska and Balog, 2024) extends CAsT-snippets (Łajewska and Balog, 2023) by including five randomly selected non-relevant passages to each question. BoolQ3L (Sulem et al., 2022) matches each selected “Yes” or “No” question of BoolQ (Clark et al., 2019) to a passage within BoolQ that has the greatest overlap with the questions in terms of nouns and verbs. These unanswerable datasets are sourced from Wikipedia, which is typically included in the training data of LLMs. Our dataset is crawled from news, allowing us to easily select news published only after the knowledge cutoff date of LLMs. Table 1 presents a comparison of these benchmark datasets.
@@ 2.2. Question Generation
Zhu et al. (2024) and Yuan et al. (2024) describe a method to synthetically generate stand-alone false premise questions by selecting a set of factual triples of (subject, relation, object) using Wikipedia data, replacing the object with a similar but incorrect one, and then filling in a template to generate a false premise question. Content-grounded (non-confusing) synthetic data generation has been studied in (Yehudai et al., 2024; Wang et al., 2023a; Lee et al., 2023). Given the grounding passages, these methods generate data entries using few-shot prompting with task-specific examples, followed by filtering to ensure data quality, faithfulness, and diversity.
@@ 2.3. Mitigation
Mitigation is commonly achieved through in-context learning  (Zhu et al., 2024; Li et al., 2024b), chain-of-thought (CoT)  (Wei et al., 2022) and fine-tuning. Experiments show that fine-tuning outperforms both in-context learning and CoT reasoning (Hu et al., 2023). CoT reasoning leads to inconsistent improvements, whereas in-context learning improves performance as the number of examples increases (Li et al., 2024b). Yuan et al. (2024) identifies approximately 1% of attention heads as responsible for confusion caused by false-premise questions and demonstrates a 20% performance gain by constraining these heads.  Kim et al. (2021) retrieves a Wikipedia page for each question and responds with the identified unverifiable premises if the page does not entail them.  Sulem et al. (2022) trains a classification model based on BERT to identify unanswerable questions.
@@ 3. Methods
We developed the ELOQ data generator to evaluate and mitigate LLM confusion when responding to out-of-scope questions for a given document. We assume that the user has limited domain knowledge, resulting in asking questions that appear relevant but are actually out of scope and cannot be answered based on the document’s content. We also assume that, due to the limitations of the retriever itself, relevant documents without an answer to the questions are possibly retrieved and ranked at the top. The best answer to this kind of question is to refuse to answer it, rather than providing a fabricated response. In this step, given a standard prompt $p$, an out-of-scope question $q$, and a relevant document $d$, the LLMr generates a response $r$. Ideally, $r$ should clarify the confusing part of the question – a process we refer to as defusion (or de-confusion) – rather than attempting to answer the question directly and risking hallucination.
3.1. Data Collection
We focus on the scenario where the domain knowledge is maintained separately from the LLM in a document database. Hence, we prefer documents that are novel to the evaluated LLMs, such as news articles published after all pretraining cutoff dates, ensuring that LLMs cannot reproduce claims or answer questions correctly by pure hallucination. Instead, they must generate responses grounded in their understanding of the news content. Therefore, using the Newscatcher (Bugara et al., 2020), we collected 200 news articles published between Jan 1, 2024, and September 30, 2024, for each topic222sport, business, science, food, politics, travel, entertainment, music, news, tech. We require each document to be concise enough to fit within the LLM prompt, along with instructions and examples, but sufficiently long to make at least 4 to 10 separate claims (i.e., a few paragraphs in length). Thus, we select news articles with more than 150 words and, for each, sequentially extract sentences from the beginning until the word count exceeds 300. Finally, we collected 2,000 news and sampled a subset for human annotation. We represent the human-annotated data as Gold and the remaining data as Silver. Table 2 presents the statistics of the dataset. Our ELOQ dataset  [^3] https://huggingface.co/datasets/zhiyuanpeng/ELOQ is publicly available.
3.2. Question Generation
3.2.1. Claim Extraction
$S_{j}=\{c_{i}\mid i\mod 3=j-1\},\quad\text{for }j=1,2,3.$
3.2.2. Hallucination Injection
The second step is to generate hallucinated claims that are similar to the factual claims generated in Section 3.2.1 but contain altered or missing information. The process follows an iterative masking and recovery approach over multiple rounds (e.g. 3). For each round, we follow steps 3 to 6 in Algorithm 1.
In practice, we find that even after repeating the hallucination injection process multiple times, some claims remain too general and can still be supported by the document. Therefore, we design a prompt-based filter  B.11 to discard these in-scope claims that are explicitly supported by $d$. This ensures that only truly out-of-scope claims remain in our dataset.
Attribute Gold Silver
# of documents 45 1,955
# of in-scope questions 103 8,949
# of out-of-scope questions 113 10,105
average words per document 280 276
average words per in-scope question 16 16
average words per out-of-scope question 16 16
Table 2. Statistics of ELOQ. Gold is sampled from crawled 2,000 news and annotated by annotators (Section 3.4).
3.2.3. Question Generation
In addition to generating out-of-scope questions, we also prompt LLMq to generate in-scope questions for comparison. The complete question generation process is outlined in Algorithm 1, and the associated prompts are in Appendix B.
Group Annotator Cohen’s Kappa Annotator Acc Group’s Agree Acc Ground Truth Acc
1 A1 0.8053 0.8917 88.89 97.37 92.31 96.97 94.91 98.23
A2 87.50 94.59
2 A3 0.7508 0.8187 93.06 97.14 98.41 96.77
A4 91.67 90.00
3 A5 0.8615 0.8333 91.67 93.75 98.51 100.00
A6 98.61 100.00
Table 3. Agreement scores and the performance of our method on the ELOQ-Gold. “Annotator” represents the annotator’s name. “Out-of-Scope” represents whether the question is out-of-scope or not. “Defusion” refers to whether an LLM’s response indicates that the question has no suitable answer based on the document.
@@ 4. Experimental Results
Table 5. Evaluation of LLMs’ accuracy in defusing out-of-scope questions from the ELOQ-Silver dataset across diverse news topics. For each LLM, the underscored value and starred () value are the minimum and maximum values, respectively, across all the topics within the same prompting method. Bold values are the maximum values for each topic across all the LLMs within the same prompting method. “entm” is the abbreviation of “entertainment”.
@@ 4.2. Out-of-Scope Detection
The primary challenge in preventing an LLM from hallucinating or generating incorrect answers to out-of-scope questions is detecting such scenarios effectively. Once we can accurately identify these cases, we can handle them easily by using a prompt and in-context examples specifically designed for out-of-scope questions. We use two methods for out-of-scope detection. First is a prompt-based method where we simply prompt the LLM to do such detection (Appendix B.3), denoted as “Direct Generate” in Figure 2. Second is a classification method where we train a binary classifier on ELOQ-Silver using the embeddings of an appended unused-token as features, denoted as “Unused-token Classification” in Figure 2.
@@ 4.2.1. Unused-token Classification
The out-of-scope detection problem can be framed simply as a binary classification task, where we determine if a given question $q$ is out-of-scope or not for a given document $d$. The binary label is $y\in\{0,1\}$ where $y=1$ indicates the question is relevant to the document and $y=0$ indicates the question is out-of-scope. The objective is to learn a classification function: $f:(q,d)\mapsto p(y=1\mid q,d)$, which estimates the likelihood of the question-document pair to be in-scope. We would like to test whether the LLM’s internal features contain enough information to classify a question-document pair effectively to simulate the effect that the LLM “realizes” the out-of-scope scenario before generating a response. To achieve this, we use the sequence representation of the entire input – denoted as $x=\text{concat}(I,q,d,)$ where $I$ is the instruction prompt provided to the LLM (e.g. “Can this document answer the question?”). We extract the sequence representation by appending a reserved unused-token to the end of the input sequence and extracting its representation from the final hidden state of the last transformer layer. The rationale behind using unused-token is that, since LLM developers reserve these tokens for potential future use in specialized training tasks, they are untrained and uninfluenced by any prior data. Thus, the unused-token does not have any predefined meanings and can be viewed as a blank buffer, aggregating interactions among all the input tokens and capturing an uninfluenced representation of the sequence. We also tried using the <eos> token, which yielded similar results.
The classifier is a two-layer MLP that takes the unused-token representation as input and is optimized against the binary label $y$. This approach is commonly known as “probing” (Ye et al., 2024), where the LLM weights are frozen, and the classifier can be viewed as an external probe to assess the information contained in the model’s internal representations. The performance reported in Figure 2 was tested on ELOQ-Gold, the human-labeled data. For more details about the training, please refer to Appendix C. We also provide a “Sentence BERT NLI” baseline where we adopt NLI style training using the off-the-shelf  [^4] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 model from HuggingFace. Specifically, we follow (Reimers and Gurevych, 2019) to extract document and question [CLS] embeddings $e_{d}$ and $e_{q}$ to use $[e_{d};e_{q};|e_{d}-e_{q}|]$ as the input features.